OpenStudy (anonymous):

Please Solve For X: x+3=2x-4

OpenStudy (anonymous):

x = 7

OpenStudy (anonymous):

\[x+3=2x-5\]\[4+3=2x-x \rightarrow x=7\]

OpenStudy (anonymous):

do you know how to get the answer?